Let A = {a, b, c}, B = {x, y} and C = {0, 1}. Find a.      C x B x A b.      B x B x B

1
Expert's answer
2021-09-21T02:56:43-0400

"C\times B\times A=\{ (0,a,x),(0,a,y),(0,b,x),(0,b,y),(0,c,x),(0,c,y)\\
(1,a,x),(1,a,y),(1,b,x),(1,b,y),(1,c,x),(1,c,y)\}"

"B\times B\times B=\{(x,x,x),(x,x,y),(x,y,x),(x,y,y),(y,x,x),(y,x,y),(y,y,x),(y,y,y)\}"